If a restaurant is 43 yards wide and 112 yards long how many square feet is it?

Answer by TutorDelphia(193)

This problem is trying to trip you up converting between yards and feet
1 yard = 3 feet
1 square yard = 9 square feet
You can either convert first using the yards to feet conversion or convert after using the square yards to square feet conversion
Converting first:
43*3*112*3=43344
Converting afterward:
43*112=4816
Convert from square yards to square feet
4816*9=43344 square feet